OpenStudy (anonymous):

Write the equation in standard form: 7 – 3(x – y) = 5x + 2 (Points : 1) 3x – 8y = 5 8x + 3y = 5 8x – 3y = 5 8x – 3y = –5

OpenStudy (johnweldon1993):

So first...distribute that -3 into the parenthesis 7 - 3(x - y) = 5x + 2 will become 7 -3x + 3y = 5x + 2 right? now combine like terms...

OpenStudy (anonymous):

im so confused....

OpenStudy (johnweldon1993):

Alright no problem...so we start with 7 - 3(x - y) = 5x + 2 the first step...is to distribute the -3 into the parenthesis....multiply the -3 by each term in the parenthesis...so -3 times x = -3x and -3 times -y = 3y so now we have 7 - 3x + 3y = 5x + 2 Now....we have 2 different 'x' numbers...we can combine them to just make 1... add 3x to both sides of the equation...so 7 - 3x + 3y = 5x + 2 +3x +3x --------------------- 7 + 3y = 8x + 2 Now...we can also combine those constant numbers as well...try subtracting 2 from both sides of the equation... 7 + 3y = 8x + 2 -2 -2 5 + 3y = 8x NOW we want both the variables on 1 side of the equation...so subtract 3y from both sides 5 + 3y = 8x - 3y -3y 5 = 8x - 3y or rewritten as 8x - 3y = 5 look familiar?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(5, 13), (1, 1), (2, –1) (–5, 13), (1, 1), (2, –1) (–5, 13), (1, 1), (2, 1) (5, 13), (–1, 1), (2, 1)

OpenStudy (anonymous):

which one???